OpenStudy (anonymous):

find the equation of the line that passes through the centers of the circles with equation x^2+y^2-6x+8y+13=0 and x^2+y^2-2y-3=0

OpenStudy (anonymous):

Find the center of each circle by converting to standard form \[x^2-6x+9+y^2+8y+16=-13+9+16\]\[(x-3)^2+(y+4)^2=12\]Center of first circle is (3,-4) Find center of second circle \[x^2+y^2-2y+1=3+1\]\[x^2+(y-1)^2=4\]Center of second circle is (0,1) (Continued...)

OpenStudy (anonymous):

Find the slope of the line\[m=\frac{-4-1}{3-0}=\frac{-5}{3}\]Use point slope and the second center since it contains a 0 coordinate \[y-1=-\frac{5}{3}(x-0)\]solve for y\[y=-\frac{5}{3}x+1\]

OpenStudy (anonymous):

how did u get the 16 and 9

OpenStudy (anonymous):

?

OpenStudy (anonymous):

do you know how to complete the square?

OpenStudy (anonymous):

oh! is that what u did there?

OpenStudy (anonymous):

I completed the square on the x-variables (the 9) and the y variables (the 16) separately Half of 6 squared is 9 x^2-6x+9 Now it is a perfect square trinomial

OpenStudy (anonymous):

thanks dear,which country u from?

OpenStudy (anonymous):

Italia, living in US right now

OpenStudy (anonymous):

school :})

OpenStudy (anonymous):

ok,good knowing u. am from nigeria,studying in canada

OpenStudy (anonymous):

we are neighbors!

OpenStudy (anonymous):

hahaha yeah

OpenStudy (anonymous):

i don't know if u can be helping me out in my assignments,especially physics and math

hero (hero):

Someone must be a Superman fan

OpenStudy (anonymous):

physics :{( math :})

OpenStudy (anonymous):

lool ok only math,am good at physics as well

OpenStudy (anonymous):

yo am going offline man catch u later u can add me on facebook (krypton krish)